Solve. x^2 + 5x + 6 = 0
Answer by ptaylor(2198)

To solve: x^2+5x+6=0
The factors of 6 are: (3)(2) and (6)(1)
In the general form of the quadratic equation (Ax^2+Bx+C=0), if A=1, then B=the sum of the factors of C. Thus, in our equation, the factors of C that add up to 5 are (3)(2) which gives us:
(x+3)(x+2)=0
x+3=0
x=-3 and
x+2=0
x=-2
We can also rewrite our equation:
x^2+5x+6=0 rewrite as:
x^2+3x+2x+6=0 which equals
x(x+3)+2(x+3)=0
(x+3)(x+2)=0 and we arrive at the same solution.
Hope this helps. Happy holidays----ptaylor